During the first 18 MiB of a 60 min trip , a car has an average speed for remaining 42 MiB so that car is having an average speed of 21 M/s for the entire trip ?

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
Total Distance voyaged = Avg speed × time 
= 21 m/s × 1 hour = 21 × 3600 = 75600 m 
Separation gone in 18 min = 18 min × 11 m/s 
= 18 × 60 × 11 = 11880 m 
Separation gone in outstanding 42 min with avg. speed of x m/s = 42 min × x m/s 
= 42 × 60 × x m/s = 2520 x m/s 
thusly 
Add up to separate = dist. in 18 min + dist. in 42 min 
75600 = 11880 + 2520 x 
2520 x = 63720 
x = 25.28 m/s
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)